Harrison Equipment of​ Denver, Colorado, purchases all of its hydraulic tubing from manufacturers in mainland China. The company has recently completed a​ corporate-wide initiative in six​ sigma/lean manufacturing. Completed oil field hydraulic system costs were reduced 4.2% over a​ one-year period, from $884,000 to $846,872. The company is now worried that all of the hydraulic tubing that goes into the systems​ (making up 20% of their total costs) will be hit by the potential revaluation of the Chinese yuan —if some in Washington get their way. How would a 11.9% revaluation of the yuan against the dollar impact total system​ costs?
Answer #1

tubing cost in dollars = total cost * 20% = 884000*0.2 = 176800

impact due to yuan revaluation = tubing cost in dollars*dollar depreciation = 176800*0.119 = 21039.2

total systems cost will increase by 21039.2 to 21039.2+846872 = 867911.2